OpenStudy (anonymous):

Look at the formula for the area of a triangle. A = 1 --- 2 bh Which is a solution for h?

OpenStudy (anonymous):

a= 1/2bh

OpenStudy (anonymous):

If you want to get h all by itself, you need to make everything that is multiplied with it 1. For instance, if I want to solve for x: \[z=3x\]Multiply both sides by 1/3\[(1/3)z=(1/3)3x\]\[z/3=1x\]\[\frac{z}{3}=x\]

OpenStudy (anonymous):

okay can you show me the steps to solvin mine

OpenStudy (anonymous):

Well I'm not going to do it for you but if you compare to mine, what will make a 1/2 become a 1?

OpenStudy (anonymous):

ooh um subtract? i suck at math excuse me if i wrong

OpenStudy (anonymous):

No. What multiplied with 1/2 makes 1?

OpenStudy (anonymous):

idk x*(1/2)=1

OpenStudy (anonymous):

Ugh. no the answer was 2. 2(1/2) = 1. So in the triangle equation multiply by 2 on both sides.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Well, do it and show me what you get. multiply by 2 on both sides of A = 1/2 b h

OpenStudy (anonymous):

multiply by 2 with 1/2?

OpenStudy (anonymous):

No. multiply by 2 on both sides of the equation.

OpenStudy (anonymous):

whats the equation

OpenStudy (anonymous):

A = bh/2

OpenStudy (anonymous):

2b * 2h?

OpenStudy (anonymous):

Oh my god dude. Sorry, this is to tough to handle. I'll just give it to you, since arithmetic is too hard. \[A = bh/2\]Multiply both sides by 2. \[2A=bh2/2\]Multiply both sides by 1/b \[2A/b = bh/b\] \[2A/b = h\]